A square and a triangle have equal perimeters. The lengths of the three sides of the triangle are 6.1 cm, 8.2 cm and 9.7 cm. What is the area of the square in square centimeters?
(A)24 (B) 25 (C) 36 (D) 48 (E) 64

Answer by stanbon(75887)

A square and a triangle have equal perimeters. The lengths of the three sides of the triangle are 6.1 cm, 8.2 cm and 9.7 cm. What is the area of the square in square centimeters?
----
Perimeter of the triangle = 6.1+8.2+9.7 = 24 cm
Therefore perimeter of the square = 24 cm
Therefore 4(side) = 24 cm
side of the square = 6 cm
Area of the square = 6^2 = 36 cm^2
=======================================
Cheers,
Stan H.
(A)24 (B) 25 (C) 36 (D) 48 (E) 64